Question 1

In recent years, Sandhill Corporation has purchased three machines. Because of frequent employee turnover in the accounting department, a different accountant was in charge of selecting the depreciation method for each machine, and various methods have been used. Information concerning the machines is summarized in the table below.
Machine Acquired Cost Residual
Value
Useful Life
(in years)
Depreciation Method
1 Jan. 1, 2018 $104,000 $12,600 5 Straight-line
2 July 1, 2019 91,250 9,150 5 Diminishing-balance
3 Nov. 1, 2019 56,434 6,570 6 Units-of-production


For the diminishing-balance method, Sandhill Corporation uses double the straight-line rate. For the units-of-production method, total machine hours are expected to be 21,680. Actual hours of use in the first 3 years were: 2019, 420; 2020, 4,110; and 2021, 5,390.

Machine 3: Units-of-production depreciation
Calculation End of Year
Year Units-of-
Production
Depreciation
Cost/Unit
Depreciation
Expense
Accumulated
Depreciation
Carrying
Amount
2019

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with correct answer

2020

Entry field with incorrect answer

Entry field with incorrect answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

2021

Entry field with incorrect answer

Entry field with incorrect answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

If Machine 2 was purchased on October 1 instead of July 1, what would be the depreciation expense for this machine in 2019? In 2020?

Depreciation expense in 2019 $

Entry field with incorrect answer

Depreciation expense in 2020 $

Entry field with incorrect answer
Answer #1

Depreciation expense per unit = ( 56434-6570)/21680 = 2.30

Machine 3: Units-of-production depreciation
Calculation End of Year
Year Units-of-
Production
Depreciation
Cost/Unit
Depreciation
Expense
Accumulated
Depreciation
Carrying
Amount
2019

Entry field with incorrect answer420

$2.30

Entry field with incorrect answer

$966

Entry field with incorrect answer

$966

Entry field with incorrect answer

55468

Entry field with correct answer

2020

4110Entry field with incorrect answer

Entry field with incorrect answer2.30

9453Entry field with correct answer

Entry field with correct answer10419

Entry field with correct answer46015

2021

5390Entry field with incorrect answer

2.30Entry field with incorrect answer

Entry field with correct answer12397

Entry field with correct answer22816

Entry field with correct answer33618

If Machine 2 was purchased on October 1 instead of July 1, what would be the depreciation expense for this machine in 2019? In 2020?
Depreciation expense in 2019 $91250*40%*3/12 = 9125

Entry field with incorrect answer

Depreciation expense in 2020 91250*40%*9/12+91250*60%*40%*3/12 = 32850

Entry field with incorrect answer